suppose the shortrun market price a competitive firm faces is birr 9 and the total cost of the firm is tc-200+q+0.02qanswer the quetions follow

a)calculate the short run equilibrium out put and profit of the firm

b)drive the mc atc and avc


1
Expert's answer
2021-12-02T20:16:17-0500

Solution:

a.). The short run equilibrium output: MR = MC

MR = 9

MC = "\frac{\partial TC} {\partial Q}" = q

9 = q

Quantity = 9

 

b.). MC = 1

ATC = "\frac{TC} {q} =\frac{200} {q} + 1 + 0.02"

AVC = VC/Q = "\frac{VC} {q} =q + \frac{0.02q} {q} = 1 + 0.02"